OpenStudy (anonymous):

P(E) =.30 P(F)= .09 P(E u F)= .36 find the conditional probability P(EIF)

OpenStudy (anonymous):

I assume P(E|F) means the probability of E or F or both.

OpenStudy (anonymous):

both

OpenStudy (anonymous):

would it be a venn diagram?

OpenStudy (anonymous):

No, we can do it this way: \[P(E \cup F) = P(E) \cdot P(E|F)\]\[\implies P(E|F) = \frac{P(E\cup F)}{P(E)}\]

OpenStudy (anonymous):

so just .36/.30?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

that was incorrect...

OpenStudy (anonymous):

Ack! sorry

OpenStudy (anonymous):

first i need to find P(E n F) and divide that by P(F)

OpenStudy (anonymous):

positive?

OpenStudy (anonymous):

No. Lemme check something

OpenStudy (anonymous):

No, that's wrong.

OpenStudy (anonymous):

Ok, I have it. \[P(F) + P(E) - P(E\cup F) = P(E\cap F)\]

OpenStudy (anonymous):

That's from the venn diagram

myininaya (myininaya):

polpak i think you wrote your first equation wrong

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

And a few other ones too. It's not my day for probability.

OpenStudy (anonymous):

It shoulda been: \[P(E\cap F) = P(F) \cdot P(E|F)\]

myininaya (myininaya):

yep i thought \cup was intersection what is it for intersection

OpenStudy (anonymous):

\cap

myininaya (myininaya):

oh

OpenStudy (anonymous):

\cup is union, (it looks like a u)

myininaya (myininaya):

\[P(E|F)=\frac{P(E \cap F)}{P(F)}\]

OpenStudy (anonymous):

Yep, then we just had to find \(P(E\cap F)\)

OpenStudy (anonymous):

And I think I had the right reasoning for that.

myininaya (myininaya):

\[P(E \cup F)=P(E)+P(F)-P(E \cap F) \] if we solve this for P( E n F) we will get what polpak has above

OpenStudy (anonymous):

Ah, good. I reasoned it a different direction, but arrived at the same place =)

myininaya (myininaya):

\[P(E \cap F)=P(E)+P(F)-P(E \cup F)\]

myininaya (myininaya):

so we need to do \[P(E|F)=\frac{P(E)+P(F)-P(E \cup F)}{P(F)}\]